A rectangle and a triangle have equal areas. The length of the rectangle is 12 inches, and its width is 8 inches. If the base of the triangle is 32 inches, what is the length, in inches, of the altitude drawn to the base?

Answer by rfer(16322)

A=lw
A=12*8
A=96in rectangle
----------------
a=1/2bh
h=a/.5b
h=96/.5(32)
h=96/16
h=6in triangle